OpenStudy (anonymous):

What is the inverse of f(x)=1/5x+3?

OpenStudy (anonymous):

\[\frac{1}{5}x + 3\]\[\frac{x}{5} + 3\]\[\frac{x}{5} + \frac{15}{5}\]\[\frac{x + 15}{5}\] Now just flip it.

OpenStudy (anonymous):